### v1

# Generic forms

Abstract.
We study forms $I=(f_{1},\ldots,f_{r})$, $\deg f_{i}=d_{i}$, in $F$ which is the free associative algebra $k\langle x_{1},\ldots,x_{n}\rangle$ or the polynomial ring $k[x_{1},\ldots,x_{n}]$, where $k$ is a field and $\deg x_{i}=1$ for all $i$. We say that $I$ has type $t=(n;d_{1},\ldots,d_{r})$ and also that $F/I$ is a $t$-presentation. For each prime field $k_{0}$ and type $t=(n;d_{1},\ldots,d_{r})$, there is a series which is minimal among all Hilbert series for $t$-presentations over fields with prime field $k_{0}$ and such a $t$-presentation is called generic if its Hilbert series coincides with the minimal one. When the field is the real or complex numbers, we show that a $t$-presentation is generic if and only if it belongs to a non-empty countable intersection $C$ of Zariski open subsets of the affine space, defined by the coefficients in the relations, such that all points in $C$ have the same Hilbert series.
In the commutative case there is a conjecture on what this minimal series is, and we give a conjecture for the generic series in the non-commutative quadratic case (building on work by Anick). We prove that if $A=k\langle x_{1},\ldots,x_{n}\rangle/(f_{1},\ldots,f_{r})$ is a generic quadratic presentation, then $\{x_{i}f_{j}\}$ either is linearly independent or generate $A_{3}$. This complements a similar theorem by Hochster-Laksov in the commutative case.
We also consider minimal quadratic presentations of universal enveloping algebras of Lie algebras and give a conjecture for the Hilbert series of a generic presentation and prove the conjecture when $r\leq n^{2}/4$ or $r\geq(n^{2}-1)/3$, where $n$ is the number of generators and $r$ is the number of relations.
Finally we show, a bit to our surprise, that the Koszul dual of a generic presentation is not generic in general. But if the relations have algebraically independent coefficients over the prime field, we prove that the Koszul dual is generic. Hereby, we give a counterexample of [25, Proposition 4.2], which states a criterion for a generic non-commutative quadratic presentation to be Koszul. We formulate and prove a correct version of the proposition.
Key words and phrases:
Generic algebra, Hilbert series, Koszul dual
1991 Mathematics Subject Classification:
13D10, 13D02, 16D50,16E30
1. Introduction
Consider sequences of forms $I=(f_{1},\ldots,f_{r})$ of certain degrees $d_{1},\ldots,d_{r}$ in the free associative $k$-algebra $F$ (the polynomial $k$-algebra $F$, respectively) in $n$ variables of degree 1. The quotient $F/I$ is called a presentation of the underlying $k$-algebra $F/I$. What should be meant by a generic presentation $F/I$? It is natural to require that, if a presentation $A=F/I$ is generic, then “almost all” sequences of the degrees $d_{1},\ldots,d_{r}$ yield presentations with the same Hilbert series as $A$, so if we pick a presentation “at random”, it will have this Hilbert series with high probability. When the field is the real or complex numbers our definition of generic given below will satisfy this requirement. In the rest of the paper $k$ will be a field and $k_{0}$ a fixed but arbitrary prime field.
Definition 1.1.
Let $F$ be the free associative algebra with $n$ generators over $k$, $k\langle x_{1},\ldots,x_{n}\rangle$, or the commutative polynomial algebra $k[x_{1},\ldots,x_{n}]$ and assume that $\deg(x_{i})=1$ for all $i$. A presentation $F/(f_{1},\ldots,f_{r})$ is said to have type $(n;d_{1},\ldots,d_{r})$ if $f_{i}$ is a form in $F$ of degree $d_{i}$ for $i=1,\ldots,r$. We assume that the degrees satisfy $2\leq d_{1}\leq d_{2}\leq\cdots\leq d_{r}$.
In case $d_{i}=2$ for all $i=1,\ldots,r$, then the type is written $(n,r)$. Let $M_{i}^{n}$ be the set of monomials of degree $i$ in $k\langle x_{1},\ldots,x_{n}\rangle$ and let $CM_{i}^{n}$ be the set of monomials of degree $i$ in $k[x_{1},\ldots,x_{n}]$.
Definition 1.2.
Let $A(z)=\sum_{n\geq 0}a_{n}z^{n}$ and $B(z)=\sum_{n\geq 0}b_{n}z^{n}$ be power series with real coefficients. We write $A(z)\geq B(z)$ if $a_{n}\geq b_{n}$ for all $n\geq 0$.
If $A$ is a presentation we denote the Hilbert series of $A$ by $A(z)$.
2. Generic forms
Some results below are certainly well-known, but we haven’t found any good references. To make the paper as self-contained as possible we choose to give them with proofs. We will use the heading “Proposition” for results which are well-known and the heading “Theorem” for results which contain some kind of novelty. We start by giving a concrete presentation of each type with minimal Hilbert series.
Proposition 2.1.
Let $t=(n;d_{1},\ldots,d_{r})$ be a type. Let $c=\{c_{m}^{i};\ m\in CM_{d_{i}}^{n},i=1,\ldots,r\}$ be a set of indeterminates. Put
$A_{t}=k_{0}(c)[x_{1},\ldots,x_{n}]/(f_{1,c},\ldots,f_{r,c})\text{ where }f_{i,c}=\sum_{m\in CM_{d_{i}}^{n}}c_{m}^{i}m.$
Then the Hilbert series $A_{t}(z)$ is minimal among all series $A(z)$ where $A$ is a commutative presentation of type t over a field $k$ with prime field $k_{0}$. Analogously, let $c=\{c_{m}^{i};\ m\in M_{d_{i}}^{n},i=1,\ldots,r\}$ be a set of indeterminates. Put
$B_{t}=k_{0}(c)\langle x_{1},\ldots,x_{n}\rangle/(f_{1,c},\ldots,f_{r,c})\text{where }f_{i,c}=\sum_{m\in M_{d_{i}}^{n}}c_{m}^{i}m.$
Then the Hilbert series $B_{t}(z)$ is minimal among all series $B(z)$ where $B$ is a non-commutative presentation of type t over a field $k$ with prime field $k_{0}$.
Proof. We prove the commutative case, the non-commutative case is similar. Let $k$ be any field with $k_{0}$ as prime field and let $A=k[x_{1},\ldots,x_{n}]/(f_{1},\ldots,f_{r})$ be any commutative presentation of type $t$. Consider the algebra $A^{\prime}=k(c)[x_{1},\ldots,x_{n}]/(f_{1,c},\ldots,f_{r,c})$ of type $t$. Then $A$ is a specialization of $A^{\prime}$. The dimension of $(f_{1},\ldots,f_{r})_{d}$ is the rank of the matrix with rows $\{m_{j}f_{i}\}$, where $m_{j}$ is a monomial of degree $d-\deg(f_{i})$ and analogously for $(f_{1,c},\ldots,f_{r,c})_{d}$. But if a subdeterminant of the latter matrix is zero, then also the corresponding subdeterminant of the former is zero. Thus $\dim(f_{1},\ldots,f_{r})_{d}\leq\dim(f_{1,c},\ldots,f_{r,c})_{d}$, and hence $\dim A_{d}\geq\dim A^{\prime}_{d}$, so $A(z)\geq A^{\prime}(z)$. Moreover $A^{\prime}=A_{t}\otimes_{k_{0}}k$, and hence $A^{\prime}(z)=A_{t}(z)$. ∎
Proposition 2.2.
Suppose the field $k$ has sufficiently high degree of transcendency over the prime field $k_{0}$, and suppose given a type $t=(n;d_{1},\ldots,d_{r})$. Then there exists a commutative presentation $A$ over $k$ (a non-commutative presentation $B$ over $k$, respectively) of type $t$ with Hilbert series equal to the minimal $A_{t}(z)$ ($B_{t}(z)$, respectively).
Proof. We prove the commutative case. Suppose $g=\{g_{m}^{i};\ m\in CM_{d_{i}}^{n},i=1,\ldots,r\}$ is a set of algebraically independent elements over $k_{0}$ in $k$. We have a map of fields $k_{0}(c)\to k$, which maps $c$ to $g$. Let $A=A_{t}\otimes_{k_{0}(c)}k$. Then $A(z)=A_{t}(z)$. ∎
Definition 2.3.
A commutative presentation $A$ over $k$ (a non-commutative presentation $B$ over $k$) of type $t$ is called generic if $A(z)=A_{t}(z)$ ($B(z)=B_{t}(z)$). Moreover, if $f_{1},\ldots,f_{r}$ are forms in $A=k[x_{1},\ldots,x_{n}]$ (in $B=k\langle x_{1},\ldots,x_{n}\rangle$, respectively) such that the presentation $A/(f_{1},\ldots,f_{r})$ (the presentation $B/(f_{1},\ldots,f_{r})$, respectively) is generic, then $f_{1},\ldots,f_{r}$ are called generic forms. A naturally graded $k$-algebra $C$ is called generic if there is a generic $t$-presentation of $C$ for some $t$ or, equivalently, a minimal presentation of $C$ is generic. The type of $C$ is the type of a minimal presentation, which is determined by the graded vector spaces $\operatorname{Tor}^{C}_{1}(k,k)$ and $\operatorname{Tor}^{C}_{2}(k,k)$.
A generic presentation of any type exists if the field $k$ has sufficiently high degree of transcendency over its prime field $k_{0}$. This follows from Proposition 2.2.
Assumption 1.
In the rest of the paper we will assume, unless otherwise stated, that the field $k$ has sufficiently high degree of transcendency over its prime field $k_{0}$, in particular $k$ may be the real or complex numbers, if $k_{0}$ is the rational numbers.
We will also study universal enveloping algebras of Lie algebras, more precisely we will study Lie superalgebras (with squares of odd elements) generated by odd generators of degree one with relations of degree two. The universal enveloping algebras of such Lie algebras are exactly the Koszul duals of commutative quadratic $k$-algebras. In the sequel we will interprete $[a,b]$ in an associative $k$-algebra as $ab+ba$ if $a,b$ are of odd degree.
Definition 2.4.
A $k$-Lie-algebra is said to have type $(n,r)$ if it is the free Lie $k$-superalgebra (with squares of odd elements) $\mathcal{L}_{k}(x_{1},\ldots,x_{n})$ on the odd elements $x_{1},\ldots,x_{n}$ modulo the Lie ideal generated by linearly independent elements $f_{1},\ldots,f_{r}$, where $f_{i}=\sum_{1\leq j\leq n}c_{i,j}x_{j}^{2}+\sum_{1\leq j<l\leq n}c_{i,j,l}([x_{j},x_{l}])$ and where $c_{i,j},c_{i,j,l}\in k$. A presentation of a $k$-Lie-algebra of type $(n,r)$ is the same but the forms $f_{1},\ldots,f_{r}$ need not be linearly independent. Similarly, if $f_{1},\ldots,f_{r}$ are interpreted as elements in $k\langle x_{1},\ldots,x_{n}\rangle$, then the presentation $k\langle x_{1},\ldots,x_{n}\rangle/(f_{1},\ldots,f_{r})$ is said to have Lie type $(n,r)$ and the underlying algebra is said to have Lie type $(n,r)$ if $f_{1},\ldots,f_{r}$ are linearly independent.
Proposition 2.5.
Let $k$ be a field with prime field $k_{0}$ and let $t=(n,r)$ be a given type. Let $c=\{c_{i,j},1\leq i\leq r,1\leq j\leq n\}\cup\{c_{i,j,l},1\leq i\leq r,1\leq j<l\leq n\}$ be a set of indeterminates. Put
$C_{t}=k_{0}(c)\langle x_{1},\ldots,x_{n}\rangle/I_{c}\text{ and }L_{t}=\mathcal{L}_{k_{0}(c)}(x_{1},\ldots,x_{n})/I_{c},$
where $I_{c}=(f_{1,c},\ldots,f_{r,c})$ and $f_{i,c}=\sum_{1\leq j\leq n}c_{i,j}x_{j}^{2}+\sum_{1\leq j<l\leq n}c_{i,j,l}[x_{j},x_{l}]$. Then $UL_{t}=C_{t}$ and the Hilbert series $C_{t}(z)$ is minimal among all series $C(z)$ where $C$ is a presentation of Lie type $(n,r)$ and $k$ has prime field $k_{0}$ and the Hilbert series $L_{t}(z)$ is minimal among all series $L(z)$ where $L$ is a $k$-Lie-algebra presentation of type $(n,r)$ and $k$ has prime field $k_{0}$. Moreover, if $C=UL$ is any $k$-algebra of Lie type $(n,r)$ and $k$ has prime field $k_{0}$, then $C(z)=C_{t}(z)$ if and only if $L(z)=L_{t}(z)$. Also, there exists a $k$-algebra $C$ of Lie type $(n,r)$ such that $C(z)=C_{t}(z)$ and there exists a $k$-Lie-algebra $L$ of type $(n,r)$ such that $L(z)=L_{t}(z)$.
Proof. Similar to the proofs of Proposition 2.1 and Proposition 2.2. The Hilbert series $UL(z)$ is obtained from $L(z)=\sum_{i=1}^{\infty}e_{i}z^{i}$ by the exponential operator
(1) $\mathrm{Exp}(L)(z)=\prod_{i=1}^{\infty}\frac{(1+z^{2i-1})^{e_{2i-1}}}{(1-z^{2i})^{e_{2i}}}.$
This operator has an inverse
(2) $\mathrm{Log}(V)(z)=\sum_{r=1}^{\infty}\frac{\mu(r)}{r}\log(V((-1)^{r+1}z^{r}))$
where $V(z)$ is a series with constant term 1, $\mu$ is the Möbius function and log is the natural logarithm, see [17]. If $L_{1}(z)=L_{2}(z)$ then the product formula (1) gives $UL_{1}(z)=UL_{2}(z)$. Also $L(z)$ is determined from $UL(z)$ by (2). Hence the claim in the theorem that $C(z)=C_{t}(z)\iff L(z)=L_{t}(z)$ follows. ∎
Definition 2.6.
A $k$-algebra presentation $C$ of a certain Lie type $t$ (a $k$-Lie-algebra presentation $L$ of a certain type $t$, respectively) is called generic if $C(z)=C_{t}(z)$ ($L(z)=L_{t}(z)$) and, similarly as before, a $k$-algebra $C$ (a $k$-Lie-algebra $L$) is called generic if its minimal presentation is generic.
As stated in Proposition 2.5 there exist generic $k$-algebras of any Lie type $(n,r)$ and generic $k$-Lie-algebras of any type $(n,r)$.
Definition 2.7.
A commutative algebra, a non-commutative algebra or a Lie algebra will be called strictly generic if all coefficients in the relations are algebraically independent over the prime field.
Let $t=(n;d_{1},\ldots,d_{r})$ be a type. A point in $k^{n^{d}}$ corresponds to a polynomial of degree $d$ in $k\langle x_{1},\ldots,x_{n}\rangle$. Thus the points in $k^{N}$, $N=\sum_{i=1}^{r}n^{d_{i}}$ are in one-to-one correspondence to non-commutative presentations of type $t$. Analogously the points in $k^{N}$, $N=\sum_{i=1}^{r}{n+d_{i}-1\choose n-1}$ are in one-to-one correspondence to commutative presentations of type $t$, and in the Lie case, the points in $k^{N}$, $n=r{n+1\choose 2}$, are in one-to-one correspondence to Lie-algebra presentations of type $t$.
Theorem 2.8.
In any of the three cases above, we consider the set of points which give generic algebras, and call them generic points. The set of generic points of the three classes in the appropriate space $k^{N}$ is a non-empty intersection of countably many Zariski-open sets. The set is Zariski-open and non-empty in the commutative case. It is also Zariski-open and non-empty for non-commutative and Lie algebras, provided there are artinian examples (nilpotent, respectively) of the corresponding type. Moreover, if $k$ is equal to the real or complex numbers then the set of points, of a given type, with Hilbert series not equal to the minimal series is a set of Lebesgue measure zero.
Proof. Let $\sum g_{j}z^{j}$ be the minimal series in any of the three classes of algebras of a certain type $t$. Let $F_{i,m}\subset k^{N}$ be the subspace of points $P$ with $\dim I_{i}<m$. Since $\dim I_{i}$ equals the rank of a large matrix, it follows that $F_{i,m}$ is closed, being the zero set of determinants. Hence the set of points $F_{i}$ with Hilbert series $\sum\lambda_{j}(P)z^{j}$ and $\lambda_{i}(P)>g_{i}$ is closed, and $\cup_{i=0}^{\infty}F_{i}$ consists of all points $P$ such that the corresponding algebra has not the minimal Hilbert series. The complement of $\cup_{i=0}^{\infty}F_{i}$ is an intersection of Zariski-open sets and it is non-empty by Proposition 2.2 and Proposition 2.5. For the commutative case, it is shown in [9, Proposition 1] that there are only finitely many possibilities for the Hilbert series of algebras of type $(n;d_{1},\ldots,d_{r})$. Since the Hilbert function eventually is a polynomial, there is a $C$, such that one of these series has smallest coefficient of $z^{i}$ for all $i\geq C$. This gives that $\cup_{i=0}^{\infty}F_{i}=\cup_{i=0}^{C}F_{i}$. We get the conclusion, that the set of algebras with minimal series is open and non-empty. Now consider the non-commutative or Lie case and suppose there are artinian (nilpotent) examples. For a proof in this case, see also [20, Appendix A, Theorem A.1]. Then there exists a $C$ such that $g_{C}=0$. Hence if $P\in F_{i}$ for some $i>C$ then $\lambda_{i}(P)>0$ and it follows that also $\lambda_{C}(P)>0$, i.e., $P\in F_{C}$. Hence $\cup_{i=0}^{\infty}F_{i}=\cup_{i=0}^{C}F_{i}$ and it follows that the complement is open. In all cases, if $k$ is equal to the real or complex numbers it is shown in [5] that an algebraic set has Lebesgue measure 0, so each $F_{i}$ has Lebesgue measure 0, and hence also $\cup_{i=0}^{\infty}F_{i}$ has measure 0. ∎
3. Minimal series
Definition 3.1.
For a power series $A(z)=\sum_{i\geq 0}a_{i}z^{i}$, let $[A(z)]=\sum_{i\geq 0}b_{i}z^{i}$, where $b_{i}=a_{i}$ if $a_{j}>0$ for all $j\leq i$ and $b_{i}=0$ otherwise.
$\sum_{i\geq 0}a_{i}z^{i}>_{lex}\sum_{i\geq 0}b_{i}z^{i}$ if for some $n$ we have $a_{i}=b_{i}$ if $i<n$ and $a_{n}>b_{n}$.
Conjecture 1.
([7]) If $A$ is a commutative generic algebra of type $(n;d_{1},\ldots,d_{r})$ then $A$ has Hilbert series $\left[\prod_{i=1}^{r}(1-z^{d_{i}})/(1-z)^{n}\right]$.
Note 1.
Let $B$ be any artinian commutative algebra of type $(n;d_{1},\ldots,d_{r})$. There are equivalent conditions for the Hilbert series to be as in the conjecture in [24] and [6].
Conjecture 1 is proved for $r\leq n$, [26], $n=2$, [7], $n=3$, [2], $r=n+1$, [27], and for many other special commutative types [10, 22, 23]. In every case, the method of proof has been the following. An example $A$ of an algebra of type $t=(n;d_{1},\ldots,d_{r})$ with Hilbert series $[\prod_{i=1}^{r}(1-z^{d_{i}})/(1-z)^{n}]$ is found. Since we have the lexicographic inequality $B(z)\geq_{lex}A(z)$ for any algebra $B$ of type $t$, we get that $A(z)$ is minimal and the conjecture is true for the type $t$.
For the non-commutative case, we need the notion of “strongly free” defined by Anick, see [1] and [3]. We choose the following definition. Let $t=(n;d_{1},\ldots,d_{r})$ be a type and let $B=k\langle x_{1},\ldots,x_{n}\rangle/(f_{1},\ldots,f_{r})$ be a presentation of type $t$. There is an exact sequence of left $B$-modules
[…]
where $S_{i}$ is a basiselement of degree $d_{i}$ for $i=1,\ldots,r$ and $T_{j}$ is a basiselement of degree 1 for $j=1,\ldots,n$. Write $f_{i}=\sum_{j=1}^{n}m_{ij}x_{j}$, where $m_{ij}$ are polynomials of degree $d_{j}-1$ for $i=1,\ldots,r$. The map $f$ is given by $f(S_{i})=\sum_{j=1}^{n}m_{ij}T_{j}$ for $i=1,\ldots,r$ and the map $g$ is given by $g(T_{j})=x_{j}$ for $j=1,\ldots,n$.
Definition 3.2.
A sequence $(f_{1},\ldots,f_{r})$ in $k\langle x_{1},\ldots,x_{n}\rangle$ where $\deg(f_{i})=d_{i}$ is called strongly free and also the quotient algebra is called strongly free if the map $f$ in the exact sequence (3) is injective.
It follows from the definition that a strongly free sequence is a minimal set of generators for the ideal they generate and that the quotient algebra $B$ has global dimension 2. The following is also a consequence of the definition.
Proposition 3.3.
A $k$-algebra $B$ of type $t=(n;d_{1},\ldots,d_{r})$ is strongly free if and only if $B(z)p_{t}(z)=1$ where $p_{t}(z)=1-nz+\sum_{i=1}^{r}z^{d_{i}}$.
Proof. Take the alternating sum of dimensions in different degrees of the exact sequence (3). ∎
In general for a $k$-algebra $B$ of type $t=(n;d_{1},\ldots,d_{r})$, the exact sequence (3) gives that $B(z)p_{t}(z)\geq 1$ and from this it is easy to prove that
(4) $B(z)\geq[p_{t}(z)^{-1}].$
Unfortunately, it is not true that the generic series of type $t$ is $[p_{t}(z)^{-1}]$. A counterexample is given in [3, Example 6.3] with $t=(2;2,7)$. However, in the quadratic case we make the following conjecture.
Conjecture 2.
If $B$ is a non-commutative generic algebra of type $(n,r)$ then $B$ has Hilbert series $[1/(1-nz+rz^{2})]$.
Compare [3, Question 6.4]. If the conjecture is true, then it follows that $B$ is artinian if and only if $r>n^{2}/4$, since the series $1/(1-nz+rz^{2})$ has negative terms if $r>n^{2}/4$ and only positive terms if $r\leq n^{2}/4$. This last fact is well known (it is part of [3, Theorem 5.1]), but we have been unable to find a proof, so as convenience for the reader we give a proof of it here.
Proposition 3.4.
The power series expansion of $1/(1-nz+rz^{2})$ has all coefficients positive if $r\leq n^{2}/4$ and has some negative coefficient if $r>n^{2}/4$.
Proof. We have $1-nz+rz^{2}=(1-\frac{n}{2}z)^{2}-(\frac{n^{2}}{4}-r)z^{2}$. Suppose $r\leq n^{2}/4$. Put $b=\sqrt{\frac{n^{2}}{4}-r}$. Then $1-nz+rz^{2}=(1-(\frac{n}{2}+b)z)(1-(\frac{n}{2}-b)z)$. We have $b\leq n/2$. If $b=n/2$, then $r=0$ and the series expansion of $1/(1-nz)$ has only positive coefficients. If $b<n/2$ the series expansions of $1/(1-(\frac{n}{2}\pm b)z)$ have only positive coefficients and hence the same is true for $1/(1-nz+rz^{2})$.
Next suppose $r>n^{2}/4$ and put $b=\sqrt{r-\frac{n^{2}}{4}}$ and $a=\frac{n}{2}+ib$. Then $1-nz+rz^{2}=(1-az)(1-\bar{a}z)$. We have $1/(1-nz+rz^{2})=A/(1-az)+\bar{A}/(1-\bar{a}z)$, where $A=a/(a-\bar{a})$. Hence $1/(1-nz+rz^{2})=\sum_{k\geq 0}(Aa^{k}+\bar{A}{\bar{a}}^{k})z^{k}$ and
$Aa^{k}+\bar{A}{\bar{a}}^{k}=2\mathrm{Re}(Aa^{k})=2\mathrm{Re}(\frac{1}{a-\bar{a}}a^{k+1})=\frac{1}{b}\mathrm{Re}(\frac{1}{i}a^{k+1})=\frac{1}{b}\mathrm{Im}(a^{k+1}).$
Since $b>0$ we have $20<\mathrm{Arg}(a)<\pi/$ 2 and hence $\pi<\mathrm{Arg}(a^{k+1})<2\pi$ for some $k$. ∎
Corollary 3.5.
Suppose Conjecture 2 is true. If $B$ is a non-commutative generic algebra of type $(n,r)$ where $r>n^{2}/4$, then the least $k$ such that $B_{k}=0$ is
$k=\left\lceil\pi/\arctan\sqrt{4r/n^{2}-1}\right\rceil-1.$
In particular, $B_{4}\neq 0$ and $B_{5}=0$ for a generic algebra of type $(n,\frac{n^{2}}{3})$, where $n$ is divisible by 3. In this case, the first non-positive coefficient in the power series expansion of $1/(1-nz+rz^{2})$ is zero.
Proof. This follows from the proof of Proposition 3.4, since $\mathrm{Arg}(a)=\arctan\sqrt{4r/n^{2}-1}$ and $\mathrm{Arg}(a)=\arctan(1/\sqrt{3})=\pi/6$ for $r=\frac{n^{2}}{3}$. Indeed,
$1/(1-nz+\frac{n^{2}}{3}z^{2})=1+nz+\frac{2n^{2}}{3}z^{2}+\frac{n^{3}}{3}z^{3}+\frac{n^{4}}{9}z^{4}+0z^{5}+\cdots.$
Theorem 3.6.
Conjecture 2 is true for $r\leq n^{2}/4$ and $r\geq n^{2}/2$ and for $n\leq 6$.
Proof. To prove the conjecture, it suffices by (4) to give examples of algebras of type $(n,r)$ with the conjectured Hilbert series. Suppose $r\leq n^{2}/4$. By Proposition 3.4 and Proposition 3.3 it is enough to prove that there is a strongly free algebra of type $(n,r)$. It is easy to see from the definition of strongly free that a subset of a strongly free set is strongly free (see also [1, Lemma 2.7]). Hence, it is enough to give an example of a strongly free algebra when $r=\lfloor n^{2}/4\rfloor$. If $r=n^{2}/4$, $n$ even, consider
$B=k\langle x_{1},\ldots,x_{n/2},y_{1},\dots,y_{n/2}\rangle/(x_{i}y_{j},1\leq i,j\leq n/2).$
Then $B$ is isomorphic as a vector space to $k\langle x_{1},\ldots,x_{n/2}\rangle\otimes k\langle y_{1},\ldots,y_{n/2}\rangle$, which has the series $1/(1-nz/2)^{2}=1/(1-nz+rz^{2})$ and it follows by Proposition 3.3 that $B$ is strongly free.
If $n=2s+1$ is odd we let the variables be $x_{1},\ldots,x_{s},y_{1},\ldots,y_{s+1}$ and define $B$ as before. Then the number of relations in $B$ is $r=s^{2}+s=\lfloor n^{2}/4\rfloor$. The series for $B$ is $B(z)=1/(1-sz)/(1-sz-z)=1/(1-nz+rz^{2})$ and hence $B$ is strongly free.
Now, suppose $r\geq n^{2}/2$. Since $1/(1-nz+rz^{2})=1+nz+(n^{2}-r)z^{2}+(n^{3}-2nr)z^{3}+\cdots$ it follows that $[1/(1-nz+rz^{2})]=1+nz+(n^{2}-r)z^{2}$ if and only if $r\geq n^{2}/2$. To prove the conjecture for $r\geq n^{2}/2$, it is thus enough to give an example of an algebra $A$ of type $(n,r)$ such that $A_{3}=0$, since then a generic algebra $A$ of type $(n,r)$ also has $A_{3}=0$ and $A$ has linearly independent relations so its series is $1+nz+(n^{2}-r)z^{2}$. It is enough to find such an example for $r=\lceil n^{2}/2\rceil$ and this is done in [3, Example 6.2] (we will come back to this example in the proof of Theorem 3.7).
A method to compute the Hilbert series of non-commutative graded algebras is given in [19]. Using this method, algebras of type $(n,r)$, where $n^{2}/4<r<n^{2}/2$, with the conjectured Hilbert series are given for $n\leq 6$ in [19], thereby proving the conjecture for $n\leq 6$. ∎
Note 2.
The best general result we know of concerning when non-commutative algebras of type $(n,r)$ are artinian is in [15] where it is shown that $r>(n^{2}+n)/4$ generic forms implies that the algebra is artinian.
In the commutative case Conjecture 1 has been proved in the first non-trivial degree $\min\{d_{i}\}+1$, [14], with generalizations in [4] and [21]. Thus, in particular, if $f_{1},\ldots,f_{r}$ are generic quadratic forms in $A=k[x_{1},\ldots,x_{n}]$, then it is proved in [14] that $\{x_{i}f_{j}\}$ are either linearly independent or generate $A_{3}$. We will prove the corresponding result in the non-commutative quadratic case.
Theorem 3.7.
Let $f_{1},\ldots,f_{r}$ be generic quadratic forms in $A=k\langle x_{,}\ldots,x_{n}\rangle$. Then $\{x_{i}f_{j}\}\cup\{f_{j}x_{i}\},1\leq i\leq n,1\leq j\leq r$, are either linearly independent or generate $A_{3}$.
Proof. For any quadratic forms $\{f_{1},\ldots,f_{r}\}$ in $A$ we have that $\{x_{i}f_{j}\}\cup\{f_{j}x_{i}\}$, $1\leq i\leq n,1\leq j\leq r$, generate a space of dimension $\leq 2nr$ with equality if and only if they are linearly independent.
Suppose $n=2s$ is even. Call the variables in $A$ $x_{1},\ldots,x_{s},y_{1},\ldots,y_{s}$. There is in [3, Example 6.2] an example with $n^{2}/2=2s^{2}$ forms $h_{1}(i,j)=x_{i}y_{j}$, $h_{2}(i,j)=x_{i}x_{j}-y_{i}y_{j}$, $1\leq i,j\leq s$, generating $A_{3}$ as a two-sided ideal. Since the number of elements in
(5) $\{x_{p}h_{k}(i,j),y_{p}h_{k}(i,j),h_{k}(i,j)x_{p},h_{k}(i,j)y_{p},1\leq p,i,j\leq s,k=1,2\}$
is $4\cdot 2s^{3}=n^{3}=\dim A_{3}$, we get that they are linearly independent. Thus also, if $f_{1},\ldots,f_{r}$ are generic quadratic forms and $r\leq n^{2}/2$, then $x_{i}f_{k},f_{k}x_{i},y_{i}f_{k},f_{k}y_{i}$ are linearly independent. Since there are $n^{2}/2$ forms that generate $A_{3}$ as a two-sided ideal, also $r\geq n^{2}/2$ generic quadratic forms generate $A_{3}$ as a two-sided ideal.
Now suppose $n=2s+1$ is odd. Call the variables $x_{1},\ldots,x_{s},y_{0},y_{1},\ldots,y_{s}$. Then $n^{2}/2=2s^{2}+2s+1/2$. We will find $2s^{2}+2s$ quadratic forms $f_{k}$ in $12s+$ 1 variables such that $x_{i}f_{k},f_{k}x_{i},y_{i}f_{k},f_{k}y_{i}$ are linearly independent, and $12s^{2}+2s+$ 1 quadratic forms $f_{k}$ in $12s+$ 1 variables such that $x_{i}f_{k},f_{k}x_{i},y_{i}f_{k},f_{k}y_{i}$, generate $A_{3}$.
We start with the $2s^{2}$ forms $h_{1}(i,j),h_{2}(i,j)$ from above of $y_{0}$-degree 0 and adjoin $2s$ new forms of $y_{0}$-degree 1. These are, for $1\leq i\leq s$,
$q_{1}(i)$ $=y_{0}x_{i}+x_{i}y_{0}+y_{i}y_{0}$
$q_{2}(i)$ $=y_{0}y_{i}+y_{i}y_{0}.$
Multiplying these $2s^{2}+2s$ forms to the right and to the left with all variables we get $(4s+2)(2s^{2}+2s)=8s^{3}+12s^{2}+4s$ expressions, which consist of three groups of $y_{0}$-degree 0,1,2, respectively. The first group are the old $8s^{3}$ expressions (5) of $y_{0}$-degree 0. The second group consist of $12s^{2}$ expressions which all have $y_{0}$-degree 1. These are, for $1\leq i,j\leq s$,
$y_{0}h_{1}(i,j),h_{1}(i,j)y_{0},y_{0}h_{2}(i,j),h_{2}(i,j)y_{0},$
$x_{i}q_{1}(j),q_{1}(j)x_{i},x_{i}q_{2}(j),q_{2}(j)x_{i},y_{i}q_{1}(j),q_{1}(j)y_{i},y_{i}q_{2}(j),q_{2}(j)y_{i}.$
Finally the third group are the following $4s$ expressions of $y_{0}$-degree 2,
$y_{0}q_{1}(i),q_{1}(i)y_{0},y_{0}q_{2}(i),q_{2}(i)y_{0}\text{ for }1\leq i\leq s.$
To show that all these forms of degree 3 are linearly independent, we can treat the three groups separately.
The first group we know are linearly independent from the even case above. To prove that the third group are linearly independent, suppose
$\sum_{i=1}^{s}\lambda_{1,i}y_{0}q_{1}(i)+\lambda_{2,i}q_{1}(i)y_{0}+\lambda_{3,i}y_{0}q_{2}(i)+\lambda_{4,i}q_{2}(i)y_{0}=0.$
The monomial $y_{0}^{2}x_{i}$ only occurs once for each $i$, hence $\lambda_{1,i}=0$ for all $i$. Looking at the monomials $x_{i}y_{0}^{2}$ and $y_{0}^{2}y_{i}$ we get in the same way that $\lambda_{2,i}=\lambda_{3,i}=0$ for all $i$. Finally also $\lambda_{4,i}=0$ for all $i$.
Next consider the second group. Suppose
$\sum_{i,j=1}^{s}$ $\lambda_{1,i,j}y_{0}h_{1}(i,j)+\lambda_{2,i,j}h_{1}(i,j)y_{0}+\lambda_{3,i,j}y_{0}h_{2}(i,j)+\lambda_{4,i,j}h_{2}(i,j)y_{0}+\lambda_{5,i,j}x_{i}q_{1}(j)+$
$\lambda_{6,i,j}q_{1}(j)x_{i}+\lambda_{7,i,j}x_{i}q_{2}(j)+\lambda_{8,i,j}q_{2}(j)x_{i}+\lambda_{9,i,j}y_{i}q_{1}(j)+$
$\lambda_{10,i,j}q_{1}(j)y_{i}+\lambda_{11,i,j}y_{i}q_{2}(j)+\lambda_{12,i,j}q_{2}(j)y_{i}$
The monomial $y_{i}x_{j}y_{0}$ occurs only in $y_{i}q_{1}(j)$, hence $\lambda_{9,i,j}=0$ for all $1\leq i,j\leq s$. In the remaining expressions the monomial $y_{0}y_{j}x_{i}$ occurs only in $q_{2}(j)x_{i}$, hence $\lambda_{8,i,j}=0$ for all $1\leq i,j\leq s$. Continuing like this, considering in turn, $x_{j}y_{0}x_{i}$, $y_{0}x_{i}x_{j}$, $x_{i}y_{0}x_{j}$, $x_{i}x_{j}y_{0}$, $y_{i}y_{j}y_{0}$, $y_{0}y_{j}y_{i}$, $y_{j}y_{0}y_{i}$, $y_{0}x_{i}y_{j}$, $x_{i}y_{0}y_{j}$, $x_{i}y_{j}y_{0}$, we get successively $\lambda_{6,i,j}=\lambda_{3,i,j}=\lambda_{5,i,j}=\lambda_{4,i,j}=\lambda_{11,i,j}=\lambda_{12,i,j}=\lambda_{10,i,j}=\lambda_{1,i,j}=\lambda_{7,i,j}=\lambda_{2,i,j}=0$ for all $1\leq i,j\leq s$.
We have thus found $2s^{2}+2s$ quadratic forms $f_{k}$ such that $x_{i}f_{k},f_{k}x_{i},y_{i}f_{k},f_{k}y_{i}$ are linearly independent. It follows that the same is true for $r$ generic quadratic forms if $r\leq 2s^{2}+2s$.
Now add $y_{0}^{2}$ to the $2s^{2}+2s$ forms above. We have generated everything in $A_{3}$ of $y_{0}$-degree 0 with the first group. The second group consists of $12s^{2}$ linearly independent elements of $y_{0}$-degree 1, and $12s^{2}$ is also the dimension of $A_{3}$ in this degree. Finally the elements in the third group together with $y_{0}^{2}$ generate everything in $A_{3}$ of $y_{0}$-degree 2 and 3.
We have thus found $12s^{2}+2s+$ 1 quadratic forms $f_{k}$ such that $\{f_{k}\}$ generate $A_{3}$ as a two-sided ideal. It follows that the same is true for $r$ generic quadratic forms if $r\geq 2s^{2}+2s+1$.
There is a natural analogue to Conjecture 2 for Lie algebras (compare [12, Conjecture 7.6] ).
Conjecture 3.
A generic Lie algebra of type $(n,r)$ has Hilbert series $\left[\mathrm{Log}(1/(1-nz+rz^{2}))\right]$, where $\mathrm{Log}$ is defined in Section 2 equation (2).
If the conjecture is true then $L$ is nilpotent if $r>n^{2}/4$, since then $1/(1-nz+rz^{2})$ has negative terms (see Proposition 3.4 and [3, Theorem 5.1]) and then also $\mathrm{Log}(1/(1-nz+rz^{2}))$ has negative terms, since if $V(z)\geq 0$ then also $\mathrm{Exp}(V)(z)\geq 0$.
We have the following analogue of Theorem 3.6.
Theorem 3.8.
Conjecture 3 is true for $r\leq n^{2}/4$ and for $r\geq(n^{2}-1)/3$. Moreover $L_{3}=0$ if and only if $r\geq(n^{2}-1)/3$ for a generic Lie algebra of type $(n,r)$.
Proof. For the first case, we proceed as in the proof of Theorem 3.6. The only difference is that we change the relations $x_{i}y_{j}$ to $[x_{i},y_{j}]$. If we choose an order with all $x_{i}$’s larger than all $y_{j}$’s, the leading monomials are $\{x_{i}y_{j}\}$, which gives that they are strongly free by [1, Theorem 3.2]. Hence $UL(z)=1/(1-nz+rz^{2})$ which gives $L(z)=\mathrm{Log}(1/(1-nz+rz^{2}))$.
We now prove the claim that a generic Lie algebra of type $(n,r)$ has $L_{3}=0$ if and only if $r\geq(n^{2}-1)/3$.
Suppose $r\geq(n^{2}-1)/3$ and suppose $f_{1},\ldots,f_{s}$ are generic quadratic forms in $k[x_{1},\ldots,x_{n}]$, where $s={n+1\choose 2}-r\leq{n+1\choose 2}-(n^{2}-1)/3={n+2\choose 3}/n$. By [14] $\{x_{i}f_{j}\}$ are linearly independent and by [16, Theorem 2.6] it follows that $L_{3}=0$ where $A^{!}=UL$ and $A=k[x_{1},\ldots,x_{n}]/(f_{1},\ldots,f_{s})$. Hence there is an example of a Lie algebra $L$ of type $(n,r)$ such that $L_{3}=0$ which implies that $L_{3}=0$ in the generic case.
Suppose on the other hand that $L_{3}=0$ for a Lie algebra of type $(n,r)$. Then by [16, Theorem 2.6] it follows that the dual relations $f_{1},\ldots,f_{s}$ satisfy $\{x_{i}f_{j}\}$ are linearly independent. But then $ns\leq{n+2\choose 3}$ from which it follows that $r={n+1\choose 2}-s\geq(n^{2}-1)/3$.
Finally we prove that the conjecture is true when $r\geq(n^{2}-1)/3$. By the above we have $L(z)=nz+sz^{2}$ for a generic Lie algebra of type $(n,r)$ with $r\geq(n^{2}-1)/3$. We have to compute the exponent $e$ such that
$1/(1-nz+rz^{2})\equiv(1+z)^{n}(1+z^{3})^{e}/(1-z^{2})^{s}\text{ modulo }z^{4}$
. A direct calculation gives $e=n(n^{2}-(3r+1))/3$ and thus
$\mathrm{Log}(1/(1-nz+rz^{2}))=nz+sz^{2}+n(n^{2}-(3r+1))/3z^{3}+\cdots$
and in particular,
$\left[\mathrm{Log}(1/(1-nz+rz^{2}))\right]=nz+sz^{2}\iff r\geq(n^{2}-1)/3$
and hence $L(z)=\left[\mathrm{Log}(1/(1-nz+rz^{2}))\right]$.
4. The Koszul dual
We may consider a quadratic non-commutative algebra $A$ with $n$ generators and $r$ linearly independent quadratic relations as a point in the Grassmanian $\mathrm{Grass}(n^{2},r)$. This is in particular useful when the Koszul dual is studied. If $U$ is a subspace of $V\otimes V$ then the Koszul dual of $A=T(V)/\langle U\rangle$ is defined as $A^{!}=T(V^{*})/\langle U^{0}\rangle$ where $U^{0}=\{f\in V^{*}\otimes V^{*};\ f(U)=0\}$. It is known that $A^{!}$ is the subalgebra of $\mathrm{Ext}_{A}(k,k)$ generated by $\mathrm{Ext}_{A}^{1}(k,k)$, see [16]. The correspondence between $A$ and $A^{!}$ may be seen as a map $\mathrm{Grass}(n^{2},r)\to\mathrm{Grass}(n^{2},n^{2}-r)$, given by $U\mapsto U^{0}$. Moreover, by performing Gaussian elimination, a basis for $U^{0}$ may be obtained using rational expressions in terms of the coefficients of a basis for $U$. Thus the map $A\to A^{!}$ is a map of projective manifolds and it is an isomorphism, since $(A^{!})^{!}=A$.
If $A$ is commutative we may consider $A$ as an element in $\mathrm{Grass}({n+1\choose 2},r)$. In this case $A^{!}$ as a quotient of $k\langle T_{1},\ldots,T_{n}\rangle$ has relations which are linear combinations of $T_{i}^{2}$ and $[T_{i},T_{j}]=T_{i}T_{j}+T_{j}T_{i}$. Hence $A^{!}$ is the enveloping algebra of a graded Lie superalgebra generated by $n$ odd elements of degree 1. The number of relations in $A^{!}$ is $n^{2}-{n\choose 2}-r={n+1\choose 2}-r$. Hence, when $A$ is commutative, the corespondence $A\leftrightarrow A^{!}$ may be seen as a map $\mathrm{Grass}({n+1\choose 2},r)\to\mathrm{Grass}({n+1\choose 2},{n+1\choose 2}-r)$.
It is reasonable to believe that if a quadratic $k$-algebra is strictly generic (i.e., has algebraically independent coefficients in the relations), then the same is true for the Koszul dual, but we have no proof. However, we can prove the following fact.
Theorem 4.1.
Let $A$ be a strictly generic quadratic $k$-algebra (commutative, non-commutative or of Lie type). Then $A^{!}$ is a generic $k$-algebra (of Lie type, non-commutative, or commutative, respectively).
Proof. We prove the non-commutative case, the other cases are similar. Let $t$ be a given type and let $B_{t}$ be the algebra introduced in Proposition 2.1. As a first step, we prove that $B_{t}^{!}$ is generic. Suppose $C$ is any $k$-algebra of the type of $B_{t}^{!}$ with prime field $k_{0}$. Then $C^{!}$ is a $k$-algebra of the type $t$ and hence $C^{!}$ is a specialization of $B_{t}$. The same specialization may be used to compute the Koszul dual of $C^{!}$, which is equal to $C$. Indeed, $(C^{!})^{!}$ is obtained by first computing $B_{t}^{!}$ and then specialize. It follows that $C(z)\geq B_{t}^{!}(z)$ and hence $B_{t}^{!}$ is generic.
Now let $A$ be any strictly generic $k$-algebra, where $k$ has prime field $k_{0}$, of type $t$. Then (with notations from Proposition 2.1) there is a map of fields $k_{0}(c)\to k$ such that $A=B_{t}\otimes_{k_{0}(c)}k$. We have $A^{!}=B_{t}^{!}\otimes_{k_{0}(c)}k$ and hence $A^{!}(z)=B_{t}^{!}(z)$ and it follows that $A^{!}$ is generic. ∎
The following theorem shows that it is essential in Theorem 4.1 to assume that $A$ is strictly generic.
Theorem 4.2.
Suppose a $k$-algebra is either
$(a)$ a generic commutative algebra or
$(b)$ a generic non-commutative algebra or
$(c)$ a generic algebra of Lie type
Then the Koszul dual is in general not generic in $(a)$ as an algebra of Lie type, in $(b)$ as a non-commutative algebra and in $(c)$ as a commutative algebra.
Proof. $(a)$
Consider the algebra of commutative type $(4,6)$,
$A=k[a,b,c,d]/(a^{2},b^{2},c^{2},ad-d^{2},ac-bd,cd).$
It is easy to see that $A_{3}=0$, hence $A$ is generic. We have
$A^{!}=k\langle a,b,c,d\rangle/([a,b],[b,c],[a,c]+[b,d],[a,d]+d^{2}).$
If $A^{!}$ is generic, then it is strongly free with series $1/(1-4z+4z^{2})$. Using e.g. the Macaulay2 [13] package “GradedLieAlgebras” [18], one can see that $L(z)=4z+6z^{2}+4z^{3}+7z^{4}+\cdots$ where $UL=A^{!}$. Developing $1/(1-4z+4z^{2})$ as an infinite product, e.g. by using (2) Section 2, we get $1/(1-4z+4z^{2})=(1+z)^{4}/(1-z^{2})^{6}(1+z^{3})^{4}/(1-z^{4})^{6}\cdots$. Hence $L(z)$ is not minimal and hence the $k$-Lie-algebra $L$ is not generic of type $(4,4)$ and then, by Proposition 2.5 also the associative algebra $A^{!}$ is not generic of Lie type $(4,4)$.
$(b)$
The algebra $A$ defined above may also be considered as a non-commutative algebra of type $(4,12)$, namely
$A=k\langle a,b,c,d\rangle/(a^{2},b^{2},c^{2},ad-d^{2},ac-bd,cd,ab-ba,ac-ca,ad-da,bc-cb,bd-db,cd-dc).$
We still have $A_{3}=0$ and $A^{!}$ is the same, so this gives the required example. We will give another example studied in [11, Lemma 4]
$B=k\langle a,b,c,d\rangle/(a^{2},b^{2},d^{2},ad,ca,cd,da,db,dc,ba,ab+ac,bc+cb,bc+c^{2}).$
The Koszul dual of $B$ is
$B^{!}=k\langle a,b,c,d\rangle/(ab-ac,bc-cb-c^{2},bd).$
Again, it is fairly easy to see that $B_{3}=0$ and hence $B$ is generic. If $B^{!}$ were generic then, by Proposition 3.4 and Theorem 3.6, its series would be $1/(1-4z+3z^{2})$. However, In [11, Lemma 4] it is proved that $B^{!}(z)=1/(1-4z+3z^{2}-z^{4})$.
$(c)$
Consider the following algebra of Lie type $(3,3)$.
$C=k\langle a,b,c\rangle/(b^{2},[a,b]-c^{2},[a,c]).$
Also in this case we can use the Macaulay2 package “GradedLieAlgebras” to get that $L_{3}=0$, where $C=UL$. Hence $L_{3}$ is generic and by Proposition 2.5 it follows that also $C$ is generic. The Koszul dual of $C$ is
$C^{!}=k[a,b,c]/(a^{2},bc,ab+c^{2}).$
Since $C$ has commutative type $(3,3)$ the generic case is a complete intersection. However, $ab$ and $ab+c^{2}$ is $\neq 0$ in $k[a,b,c]/(a^{2},bc)$ but $ab(ab+c^{2})=0$ in $k[a,b,c]/(a^{2},bc)$. This example, under the name 3b, was studied in [8]. ∎
Note 3.
In the proof of [25, Proposition 4.2] it is used that the Koszul dual of a generic quadratic algebra is generic. The definition of generic in [25] is that the algebra should belong to a countable intersection $C$ of non-empty Zariski open subsets of the appropriate affine space. We think that one has to add the assumption that the Hilbert series is constant in $C$, so by Theorem 4.3 below it coincides with our definition. But even with this assumption it is not true that the Koszul dual of a generic algebra is generic, by Theorem 4.2. Indeed, the example provided in the proof of b) in Theorem 4.2 gives a counterexample also to [25, Proposition 4.2], since $A$ satisfies their condition $s\leq m^{2}/4$, $A$ is not Koszul since $A^{!}$ is not Koszul and $A$ is generic since $A_{3}=0$. We give in Theorem 4.4 below a correct version of [25, Proposition 4.2].
Theorem 4.3.
A commutative presentation of type $t$ is generic if and only if it belongs to a non-empty Zariski-open subset of the affine space of commutative presentations of type $t$, whose points have constant Hilbert series. A non-commutative or Lie presentation over the real or complex numbers of type $t$ is generic if and only if it belongs to a non-empty countable intersection of Zariski-open subset of the affine space of non-commutative or Lie presentations of type $t$, whose points have constant Hilbert series.
Proof. If the presentation is generic, the result follows from Theorem 2.8. Suppose $A$ is a commutative presentation of type $t$, which belongs to a non-empty Zariski-open subset $O$ of the affine space of commutative presentations of type $t$ with the property that all points in $O$ have the same Hilbert series.. By Theorem 2.8, the set $G$ of all generic points is open and non-empty. Since the field $k$ is infinite, $O\cap G$ is non-empty and from this it follows that $A$ is generic. Suppose now that $A$ is a non-commutative or Lie presentation of type $t$ over the real or complex numbers, which belongs to a non-empty countable intersection $O$ of Zariski-open subsets of the affine space of non-commutative or Lie presentations of type $t$ with the property that all points in $O$ have the same Hilbert series. By Theorem 2.8, the set $G$ of all generic points is a non-empty countable intersection of open sets. By [5] it follows that $G\cap O$ is non-empty since its complement is a countable union of sets of Lebesgue measure zero. We may conclude as before that $A$ is generic. ∎
Theorem 4.4.
A strictly generic non-commutative algebra of type $(n,r)$ is Koszul if and only if $r\leq n^{2}/4$ or $r\geq 3n^{2}/4$.
Proof. Suppose first $A$ is a generic non-commutative algebra of type $(n,r)$. If $r\leq n^{2}/4$ then, by Theorem 3.6, Proposition 3.4 and Proposition 3.3, $A$ is strongly free and thus has global dimension $\leq 2$. Hence $A$ is Koszul, since $\operatorname{Tor}^{A}(k,k)$ lies on the diagonal. If $n^{2}/2\leq r<3n^{2}/4$, then by [3, Example 6.2], $A_{3}=0$ and if $A$ is Koszul then $A^{!}$ has Hilbert series $1/(1-nz+sz^{2})$, where $s=\dim(A_{2})=n^{2}-r>n^{2}/4$. However, by Proposition 3.4, the series $1/(1-nz+sz^{2})$ has some negative coefficient. Thus we have proved that if $A$ is generic of type $(n,r)$, then $A$ is Koszul if $r\leq n^{2}/4$ and not Koszul if $n^{2}/2\leq r<3n^{2}/4$.
Suppose now that $A$ is strictly generic of type $(n,r)$ and $n^{2}/4<r\leq n^{2}/2$. Then by Theorem 4.1 $A^{!}$ is generic and its type is $(n,s)$ where $n^{2}/2\leq s<3n^{2}/4$. By the above we conclude that $A^{!}$ is not Koszul and hence also $A$ is not Koszul. Finally, suppose $A$ is strictly generic of type $(n,r)$ where $r\geq 3n^{2}/4$. Then by Theorem 4.1 $A^{!}$ is generic of type $(n,s)$ where $s\leq n^{2}/4$. Hence by the above $A^{!}$ is Koszul and hence also $A$ is Koszul. ∎
There is no conflict of interest. All data is created by the author. No funding was received.